2. Fill in the missing information. (Round the "Turnover" and "ROI" answers to 2 decimal places.)

Company
A B C
Sales $5,061,000 $1,330,000
Net operating income $708,540 $186,200
Average operating assets $2,410,000 $3,320,000
Margin % % 8 %
Turnover 2.00
Return on investment (ROI) % 7.00 % %
Answer #1

Margin=Net operating income/Sales

Turnover=Sales/Average operating assets

ROI=Net operating income/Average operating assets

A B C
Sales 5,061,000 1,330,000 (3,320,000*2)=$6,640,000
Net operating income 708540 186200 (6,640,000*8%)=$531200
Average operating assets 2,410,000 (186200/0.07)=$2,660,000 3,320,000
Margin (708540/5,061,000)=14% (186200/1,330,000)=14% 8%
Turnover (5,061,000/2,410,000)=2.1 (1,330,000/2,660,000)=0.5 2
ROI (708540/2,410,000)=29.4% 7% (531200/3,320,000)=16%